Use the given information to find sin2θ, cos2θ, tan2θ, for sinθ=12/13.
Answer by Alan3354(69443)

sin(t) = 12/13
cos(t) = sqrt(1 - sin^2(t)) = 5/13
------------
sin(2t) = 2sin(t)cos(t)
----------------
cos(2t) = sqrt(1 - sin^2(2t))
tan(2t) = sin(2t)/cos(2t)